OpenStudy (anonymous):

What is the transverse axis and conjugate axis of the hyperbola y^2/49-x^2/36=1 ? The transverse axis is 12 and the conjugate axis is 14. The transverse axis is 14 and the conjugate axis is 12. The transverse axis is 36 and the conjugate axis is 49. The transverse axis is 49 and the conjugate axis is 36.

OpenStudy (anonymous):

how do u know?

OpenStudy (anonymous):

The transverse axis is y=0 and the conjugate axis is x=0. That's what I think with transverse axis and conjugate axis. Maybe I'm wrong, you better wait for other users.

OpenStudy (anonymous):

\[\mathrm{\frac{y^2}{49}-\frac{x^2}{36}=1\\\frac{y^2}{7^2}-\frac{x^2}{6^2}=1}\] a=7 and b=6

OpenStudy (anonymous):

The transverse axis is 12 and the conjugate axis is 14.

OpenStudy (anonymous):

I think is just a definition (2a transverse axis and 2b conjugate axis). With "transverse axis" I understand a line (the line which contained the 2 foci). I think there's a different definition in english than in spanish, that's why i said y=0 and x=0.